Joe has applied to two different universities. The acceptancerate for university A is 25% the acceptance rate for university Bis 40% What is the probability that Joe will be accepted by atleast one of these two universities?

the answer should be 0.55 but please show why and how
Answer #1

P(Joe will be accepted in University A) = 0.25

P(Joe will not be accepted in University A) = 1 - 0.25 = 0.75

P(Joe will be accepted in University B) = 0.40

P(Joe will not be accepted in University B) = 1 - 0.40 = 0.60

P(Joe will be accepted by at least one of these two universities) = 1 - P(Joe will not be accepted by both Universities)

= 1 - 0.75x0.60

= 0.55